OpenStudy (anonymous):

What is the equation of the parabola with a directrix at y = -4 and focus (0, 4)?

OpenStudy (anonymous):

A. x=-1/16y^2 B. y=-1/16x^2 C. y=1/16x^2 D. x=1/16y^2

OpenStudy (anonymous):

So what is your vertex?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):